OpenStudy (anonymous):

What experssion is equivalent to x^6 x^2

OpenStudy (anonymous):

x^(6+2)

OpenStudy (kainui):

What is an exponent? It's just a shortcut to writing out this: \[\LARGE x^3=x*x*x\] So if you have two things multiplied together: \[x^3*x^2=(x*x*x)*(x*x)=x*x*x*x*x = x^5\] see how you can figure out the rule?

OpenStudy (anonymous):

This is an expression with the same base so we simply add the power!! therefore it will be x^8